OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now? Numerical Answers Expected!

OpenStudy (anonymous):

C = coach; T = Tim C = T + 10; C-3 = 2(T-3), so C = 2T - 6 + 3 = 2t - 3 2t - 3 = t + 10 t = 13, Tim's age now, so coach is 23; three years ago they were 10 & 20

OpenStudy (anonymous):

let age of Tim is x yrs Jonson age is x+10 three yrs ago age of Tim=x-3 and age of coach=x+7 but x+7=2(x-3) x+7=2x-6 x=13 yrs:)